Solve the system by addition.
5x – 3y = 13
4x – 3y = 11

Answer by malakumar_kos@yahoo.com(315)


Solve the system by addition.
5x – 3y = 13
4x – 3y = 11
5x- 3y = 13.........(1)
4x - 3y = 11........(2)
multiply (1) by -1 and add to (2)
-5x+3y = -13
4x-3y = 11
-x = -2 (cancelling 3y &-3y)
x = 2
substituting for x in (1)
5*2 - 3y = 13
10 - 3y =13
10-13 = 3y
3y = -3 or y =-1
The required solution is x = 2 and y = -1